OpenStudy (anonymous):

A recipe for punch says to use 1 1/6 cups of lemonade for each cup of ginger ale. If you use 1 2/5 cups of ginger ale. How much lemonade should you use? Do I still multiply or do I divide? Not sure

OpenStudy (anonymous):

in the begining you have the ratio 7/6 : 1 ( 7/6 lemonade to 1 ginger) x : 1 2/5 or x: 7/5

OpenStudy (anonymous):

if you multiply the 7/6 with 7/5 you will get the right amount for lemonade

OpenStudy (anonymous):

I got 1 19/30...Not sure if that's correct

OpenStudy (michele_laino):

I think your answer is: \[\left( 1+\frac{ 1 }{ 6 } \right)*\frac{ 5 }{ 7 }=\frac{ 7 }{ 6 }*\frac{ 5 }{ 7 }=\frac{ 5 }{ 6 }\]

OpenStudy (michele_laino):

because you have to multiply (1 1/6) times the inverse of (1 2/5)

OpenStudy (anonymous):

ok, thank you!

OpenStudy (michele_laino):

thank you!